# Thermodynamic analysis of shift-symmetric black-hole spacetimes in Horndeski gravity

### v1

@@ Abstract
In this study, we investigate the thermodynamic properties of shift-symmetric Horndeski and beyond Horndeski theories (theories in which only derivatives of the scalar field appear in the action). Utilizing Euclidean methods, we first analyze two specific cases that serve as foundational examples for the broader framework. We derive the general expression for entropy variation within this setting, for homogeneous spacetimes, demonstrating that both parity-preserving and parity-violating terms contribute to the entropy formula. Given the functional form of the coupling terms, the black hole entropy can be determined by integrating this relation with respect to the event horizon. Our general result is consistent with previously reported special cases in the literature. Notably, for particular constraints on the coupling functions, the entropy identically vanishes. Furthermore, we establish that homogeneous spacetimes within shift- and parity-symmetric beyond-Horndeski theories universally obey Bekenstein’s Area Law for entropy, irrespective of the specific form of the coupling functions.
@@ I Introduction
A particularly important subclass of Horndeski gravity is shift-symmetric Horndeski theory, where the scalar field enters the action only through its derivatives. This additional symmetry has profound theoretical and practical implications. Many shift-symmetric scalar-tensor models naturally arise from higher-dimensional theories, such as compactifications of Lovelock gravity [19, 20, 21, 22, 23, 24], where the shift symmetry emerges as a remnant of higher-dimensional diffeomorphism invariance. Similarly, low-energy limits of Einstein-dilaton-Gauss-Bonnet gravity often exhibit shift symmetry [25, 26], reinforcing its significance in effective field theory descriptions of modified gravity. Beyond its theoretical motivation, shift symmetry plays a crucial role in simplifying the field equations, often enabling the derivation of exact analytical solutions that would otherwise be inaccessible [27]. This has been particularly important for the discovery of stealth black hole solutions, where the scalar field has a non-trivial configuration while leaving the metric identical to GR solutions [16, 28]. Furthermore, shift symmetry is essential for constructing solutions with primary scalar hair [29, 30, 31], providing a mechanism through which black holes can support an independent scalar charge.
@@
Horndeski theory has been extensively utilized in the study of exact gravitational solutions, leading to the discovery of black holes, wormholes, and solitonic configurations that deviate from their counterparts in GR. Black hole solutions in this framework have been of particular interest, as they offer testable deviations from the Kerr [37, 38, 39, 40, 41] and Schwarzschild metrics [42, 43, 44, 45, 46, 47, 48, 49, 50, 48, 51, 52, 53, 54, 55, 56, 57, 58] potentially revealing new gravitational effects in the strong-field regime. Traditionally, black holes in scalar-tensor theories were thought to obey strict no-hair theorems, which prohibit the existence of independent scalar charges. However, recent studies have demonstrated that within Horndeski gravity, black hole solutions featuring primary scalar hair can indeed exist [29, 30, 31, 59]. These configurations are particularly intriguing, as the scalar field is not merely a secondary effect of the black hole’s mass or charge but instead constitutes an independent characteristic of the solution. Such discoveries challenge the traditional understanding of black hole structure and open new avenues for exploring deviations from GR. In particular, in the same shift-symmetric theories that support the primary hair, a large number of novel regular solutions, i.e. solutions without an initial singularity, have been found [29, 31]. In addition to black holes, Horndeski gravity has also been employed in constructing wormhole solutions, where the scalar field plays a crucial role in supporting the exotic geometry [27, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71], and solitonic solutions, which describe localized, self-stabilizing field configurations in curved spacetime [27, 29, 31, 72, 73]. The emergence of these solutions highlights the richness of Horndeski theory and its potential to describe novel gravitational phenomena beyond standard GR predictions.
@@
In scalar-tensor theories, and particularly in Horndeski gravity, black hole thermodynamics plays a fundamental role in understanding the interplay between additional scalar degrees of freedom and gravitational dynamics. Since scalar-tensor theories introduce non-minimal couplings between the scalar field and curvature, they generally modify the standard thermodynamic properties of black holes. The entropy, temperature, and free energy can acquire novel corrections, leading to distinct thermodynamic behavior compared to GR solutions. A characteristic example concerns Bekenstein’s area law for black hole entropy [74, 75], which states that the entropy of a black hole is proportional to the area of the event horizon. This observation was made possible due to Hawking’s proof that the area of the event horizon never decreases [82]. In particular, it has been showed so far that the entropy of black holes in scalar-tensor theories may not follow the area law and instead obey a modified area law due to the fact that Newton’s constant has been promoted to a field that depends on the scalar degree of freedom introduced by modified gravity [83]. Particular examples of this behavior include black holes with conformally coupled scalar fields [84, 47, 48, 85, 86], black holes in $f(R)$ gravity [87, 88] and Brans-Dicke gravity [89], a phenomenon which however is absent in the case of minimally coupled scalar fields [90, 91, 92, 93, 49, 94, 45, 44]. Such behavior has been observed in general in modified gravity [95, 96, 97]. Investigating these modifications is crucial for determining whether Horndeski gravity respects established thermodynamic principles.
@@
The thermodynamics of black holes in Horndeski gravity has been an active area of research, with several works addressing this topic from different perspectives. Early studies focused on specific black hole solutions, extracting their entropy and temperature within particular subclasses of Horndeski theory [98, 99, 100, 101]. More recently, efforts have been made to understand black hole thermodynamics in the most general setting possible. Notable examples include works such as [102], which explored the thermodynamic properties of black holes in Horndeski gravity using the Iyer-Wald formalism [103, 104] providing a detailed study of black hole thermodynamics in general Horndeski theories. However, given the vast complexity of Horndeski and beyond-Horndeski theories, further investigation is needed to develop a systematic method for extracting thermodynamic properties in a broad class of solutions.
@@
In this work, we develop a general approach to extract black hole thermodynamics in shift-symmetric Horndeski and beyond-Horndeski theories by applying the Euclidean method. Our method allows us to compute the thermodynamic quantities directly from the action, providing a systematic framework for analyzing black hole solutions in these extended theories. By implementing this formalism, we derive the entropy, the temperature, and the mass for a wide class of black hole spacetimes, including those with non-trivial scalar hair. This approach not only unifies existing thermodynamic analyses but also provides new insights into the role of shift symmetry and beyond-Horndeski modifications in black hole physics. Our results contribute to the broader effort of understanding black hole thermodynamics in modified gravity and serve as a stepping stone for future studies on the thermodynamic stability and phase transitions of black holes in scalar-tensor theories. In particular, we have derived a general expression for the variation of the entropy in any shift-symmetric beyond-Horndeski theory. The black hole entropy can be determined by simply integrating this relation with respect to the event horizon radius, ensuring that it vanishes in the absence of the black hole (see section IV) This general result, applies to any homogeneous metric and has been tested by us in various well-known black-hole solutions. As is already pointed out in [96, 98, 105, 102], the Wald entropy formula [103] may not be applicable in the Horndeski framework, due to the covariant derivatives of the scalar field in the action, while the Wald method may lead to inconsistent results [98]. Our approach relies on the fact that we obtain the exact form of the boundary term that needs to be added in order to have a well defined variational procedure.
@@ II Theoretical Framework
Here, the notation is defined as follows: $\phi_{\mu}=\partial_{\mu}\phi,~{}\phi_{\mu\nu}=\nabla_{\mu}\partial_{\nu}\phi$ and subscripts denote derivative with respect to the argument. GR is included in the $\mathcal{L}_{4}$ term, when $G_{4}$ contains a constant, $X-$independent term. The beyond Horndeski terms parametrized by $F_{4}$ and $F_{5}$ are not independent. They are related so as to evade the appearance of a ghost degree of freedom [106]. This relation reads
@@ III.1.1 The theory and the solution
For the coupling functions
$G_{2}=-\varepsilon\mu X^{2}~{},~{}G_{4}=1-\frac{1}{2}X(\delta\zeta-\alpha)-\frac{1}{2}\delta\mu X^{2}~{},~{}F_{4}=\frac{3\delta\mu}{8}-\frac{\alpha-\delta\zeta}{8X}~{}.$ (12)
we will solve the field equations in order to obtain the spacetime element as well as the form of the scalar field, which we will assume that inherits the spacetime symmetries and therefore is a function of the radial coordinate. Here $\alpha,\delta,\varepsilon,\zeta,\mu$ are constants of the theory. Note, that only derivatives of the scalar field appear in the action and hence we are free to substitute $\phi^{\prime}(r)$ with $X(r)$, using relation (10). Assuming the line element
@@
with $m$ being an integration constant related to the ADM mass of the black hole $\lim_{r\to\infty}(r-rf(r))/2=m$ Note that the form of the kinetic energy does not contain the integration constant $m$ and that its behavior is solely determined by the constants of the theory. When $\alpha=\delta\zeta$, $X=0$ and $f(r)=1-2m/r$, hence we go back to the Schwarzschild solution.
@@ III.1.2 Euclidean Action
where the subscript $E$ denotes Euclidean quantities, and $\psi$ collectively represents the matter fields. The spacetime element has an Euclidean signature (+,+,+,+)(+,+,+,+)( + , + , + , + ), and we can transition to this signature by performing a Wick rotation $t\to i\tau$. If we are considering a black hole spacetime, then to avoid the conical singularity at the horizon of the black hole $r_{h}$, the time coordinate $\tau$ must be periodic, with a period given by [108, 49]
@@
By imposing the saddle point approximation, we may consider that for the black hole spacetime that will concern us in this work, it is sufficient to consider the contribution of the action of the theory evaluated on the classical solution when the field equations hold: $\delta\mathcal{I}_{E}=0$ Consequently, we need to calculate the on-shell action of our theory, ensuring it attains a true extremum within the class of fields considered. From this, we can relate the Euclidean action to the free energy of the system via
@@
[…]
where $P$ is the corresponding conjugate momentum of the scalar field, $\Pi$ is the momentum of the metric, $N$ is the lapse function, $N^{i}$ is the shift vector, and $H,~{}H^{i}$ are the Hamiltonian densities. $\mathcal{B}$ denotes the boundary terms required to ensure a well-defined variational problem. For a very pedagogical explanation on why the boundary terms should be included, we refer the reader to [110]. Since both the scalar field and the spacetime metric are time independent, while the spacetime is non-rotating we can consider the reduced action principle
[…]
@@
[…]
@@ III.1.3 Thermodynamic Quantities from Boundary Terms
$-\alpha+\delta\zeta+2\mu X(r)\left(\delta+\varepsilon r^{2}\right)~{},$ (26)
@@
[…]
@@
[…]
@@
Figure 1: The temperature as a function of the horizon radius. Here we have set $\delta\zeta/\alpha=4$,$\delta/\varepsilon=2$ for three different ratios of $\mu/\zeta$.
@@ III.2 Parity Violating Theories
In this subsection we will discuss the thermodynamics of a black hole solution arising when the parity breaking term of $F_{5}(X)$ is present. An exact solution has already been obtained in [27]. For $F_{4}(X)=0=G_{5}(X)$ and the coupling functions
@@
[…]
@@
However, at the event horizon of the black hole we have $f(r_{h})=0$ and hence no boundary term is generated at the event horizon of the black hole, and by comparing with the previous discussion this will imply that these solutions have a vanishing entropy. Consequently, a well-defined variational principle implies

### v2

@@ Abstract
In this study, we investigate the thermodynamic properties of shift-symmetric Horndeski and beyond Horndeski theories (theories in which only derivatives of the scalar field appear in the action). Utilizing Euclidean methods, we first analyze two specific cases that serve as foundational examples for the broader framework. We derive the general expression for entropy variation within this setting, for homogeneous spacetimes (with $g_{tt}=1/g_{rr}$), demonstrating that both parity-preserving and parity-violating terms contribute to the entropy formula. Given the functional form of the coupling terms, the black hole entropy can be determined by integrating this relation with respect to the event horizon. Our general result is consistent with previously reported special cases in the literature. Notably, for particular constraints on the coupling functions, the entropy identically vanishes. Furthermore, we establish that homogeneous spacetimes within shift- and parity-symmetric beyond-Horndeski theories universally obey Bekenstein’s Area Law for entropy, irrespective of the specific form of the coupling functions.
@@ I Introduction
A particularly important subclass of Horndeski gravity is shift-symmetric Horndeski theory, where the scalar field enters the action only through its derivatives. This additional symmetry has profound theoretical and practical implications. Many shift-symmetric scalar-tensor models naturally arise from higher-dimensional theories, such as compactifications of Lovelock gravity [19, 20, 21, 22, 23, 24], where the shift symmetry emerges as a remnant of higher-dimensional diffeomorphism invariance. Similarly, low-energy limits of Einstein-dilaton-Gauss-Bonnet gravity often exhibit shift symmetry [25, 26], reinforcing its significance in effective field theory descriptions of modified gravity. Beyond its theoretical motivation, shift symmetry plays a crucial role in simplifying the field equations, often enabling the derivation of exact analytical solutions that would otherwise be inaccessible [27]. This has been particularly important for the discovery of stealth black hole solutions, where the scalar field has a non-trivial configuration while leaving the metric identical to GR solutions [16, 28]. Furthermore, shift symmetry is essential for constructing solutions with primary scalar hair [29, 30, 31], providing a mechanism through which black holes can support an independent scalar charge, while, besides its application in black-hole physics it has also been used in order to alleviate cosmological tensions [32].
@@
Horndeski theory has been extensively utilized in the study of exact gravitational solutions, leading to the discovery of black holes, wormholes, and solitonic configurations that deviate from their counterparts in GR. Black hole solutions in this framework have been of particular interest, as they offer testable deviations from the Kerr [38, 39, 40, 41, 42] and Schwarzschild metrics [43, 44, 45, 46, 47, 48, 49, 50, 51, 49, 52, 53, 54, 55, 56, 57, 58, 59] potentially revealing new gravitational effects in the strong-field regime. Traditionally, black holes in scalar-tensor theories were thought to obey strict no-hair theorems, which prohibit the existence of independent scalar charges. However, recent studies have demonstrated that within Horndeski gravity, black hole solutions featuring primary scalar hair can indeed exist [29, 30, 31, 60, 61]. These configurations are particularly intriguing, as the scalar field is not merely a secondary effect of the black hole’s mass or charge but instead constitutes an independent characteristic of the solution. Such discoveries challenge the traditional understanding of black hole structure and open new avenues for exploring deviations from GR. In particular, in the same shift-symmetric theories that support the primary hair, a large number of novel regular solutions, i.e. solutions without an initial singularity, have been found [29, 31]. In addition to black holes, Horndeski gravity has also been employed in constructing wormhole solutions, where the scalar field plays a crucial role in supporting the exotic geometry [27, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73], and solitonic solutions, which describe localized, self-stabilizing field configurations in curved spacetime [27, 29, 31, 74, 75]. The emergence of these solutions highlights the richness of Horndeski theory and its potential to describe novel gravitational phenomena beyond standard GR predictions.
@@
In scalar-tensor theories, and particularly in Horndeski gravity, black hole thermodynamics plays a fundamental role in understanding the interplay between additional scalar degrees of freedom and gravitational dynamics. Since scalar-tensor theories introduce non-minimal couplings between the scalar field and curvature, they generally modify the standard thermodynamic properties of black holes. The entropy, temperature, and free energy can acquire novel corrections, leading to distinct thermodynamic behavior compared to GR solutions. A characteristic example concerns Bekenstein’s area law for black hole entropy [76, 77], which states that the entropy of a black hole is proportional to the area of the event horizon. This observation was made possible due to Hawking’s proof that the area of the event horizon never decreases [84]. In particular, it has been showed so far that the entropy of black holes in scalar-tensor theories may not follow the area law and instead obey a modified area law due to the fact that Newton’s constant has been promoted to a field that depends on the scalar degree of freedom introduced by modified gravity [85]. Particular examples of this behavior include black holes with conformally coupled scalar fields [86, 48, 49, 87, 88, 89, 90], black holes in $f(R)$ gravity [91, 92] and Brans-Dicke gravity [93], a phenomenon which however is absent in the case of minimally coupled scalar fields [94, 95, 96, 97, 50, 98, 99, 46, 45, 100]. Such behavior has been observed in general in modified gravity [101, 102, 103]. Investigating these modifications is crucial for determining whether Horndeski gravity respects established thermodynamic principles.
@@
The thermodynamics of black holes in Horndeski gravity has been an active area of research, with several works addressing this topic from different perspectives. Early studies focused on specific black hole solutions, extracting their entropy and temperature within particular subclasses of Horndeski theory [104, 105, 106, 107]. The thermodynamics of regular black holes arising from an infinite tower of curvature corrections was studied more recently in [108] (for interesting studies regarding this matter and regular black holes arising in the context of pure gravity theories we refer to [109, 110, 111, 112, 113]). Efforts have also been made to understand black hole thermodynamics in the most general setting possible. Notable examples include works such as [114], which explored the thermodynamic properties of black holes in Horndeski gravity using the Iyer-Wald formalism [115, 116] providing a detailed study of black hole thermodynamics in general Horndeski theories. However, given the vast complexity of Horndeski and beyond-Horndeski theories, further investigation is needed to develop a systematic method for extracting thermodynamic properties in a broad class of solutions.
@@
In this work, we develop a general approach to extract black hole thermodynamics in shift-symmetric Horndeski and beyond-Horndeski theories by applying the Euclidean method. Our method allows us to compute the thermodynamic quantities directly from the action, providing a systematic framework for analyzing black hole solutions in these extended theories. By implementing this formalism, we derive the entropy, the temperature, and the mass for a wide class of black hole spacetimes, including those with non-trivial scalar hair. This approach not only unifies existing thermodynamic analyses but also provides new insights into the role of shift symmetry and beyond-Horndeski modifications in black hole physics. Our results contribute to the broader effort of understanding black hole thermodynamics in modified gravity and serve as a stepping stone for future studies on the thermodynamic stability and phase transitions of black holes in scalar-tensor theories. In particular, we have derived a general expression for the variation of the entropy in any shift-symmetric beyond-Horndeski theory. The black hole entropy can be determined by simply integrating this relation with respect to the event horizon radius, ensuring that it vanishes in the absence of the black hole (see section IV) This general result, applies to any homogeneous metric and has been tested by us in various well-known black-hole solutions. As is already pointed out in [102, 104, 117, 114], the Wald entropy formula [115] may not be applicable in the Horndeski framework, due to the covariant derivatives of the scalar field in the action, while the Wald method may lead to inconsistent results [104]. The Wald entropy formula has been tested many times in the literature so far and in general yields results that agree with Euclidean methods. However, as is pointed out in [118, 102] in several theories an inconsistency may arise. In particular, Ref. [102] showed that the ambiguity is harmless for stationary black holes with a regular Killing horizon, but in scalar–tensor theories the regularity assumption can fail, e.g. due to divergences of scalar derivatives at the horizon [118], and the Wald formula may then become ill-defined. The theories are equivalent (yield the same equations of motion) up to boundary terms and the Wald formula relies on the derivative of the Lagrangian with respect to the Riemann tensor, while treating the metric as an independent quantity. It is not clear whether taking this derivative of a theory, before or after integration by parts will give the same result for the Wald entropy and this seems to be the reason for inconsistent results. The Euclidean method evades these problems, as is also discussed in [117]. The sole requirement of the Euclidean procedure is the attainment of a well-defined variational procedure by adding the appropriate boundary term and our approach relies on the fact that we obtain the exact form of the boundary term that needs to be added in order to achive this.
@@ II Theoretical Framework
Here, the notation is defined as follows: $\phi_{\mu}=\partial_{\mu}\phi,~\phi_{\mu\nu}=\nabla_{\mu}\partial_{\nu}\phi$ and subscripts denote derivative with respect to the argument: $G_{iX}=\partial_{X}G_{i}$. GR is included in the $\mathcal{L}_{4}$ term, when $G_{4}$ contains a constant, $X$-independent term. The beyond Horndeski terms parametrized by $F_{4}$ and $F_{5}$ are not independent. They are related so as to evade the appearance of a ghost degree of freedom [119]. This relation reads
@@ III.1.1 The theory and the solution
The coupling functions
$G_{2}=-\varepsilon\mu X^{2}~,~G_{4}=1-\frac{1}{2}X(\delta\zeta-\alpha)-\frac{1}{2}\delta\mu X^{2}~,~F_{4}=\frac{3\delta\mu}{8}-\frac{\alpha-\delta\zeta}{8X}~,$ (12)
have been used in [27] for the derivation of analytical solutions. We can solve the field equations in order to obtain the spacetime element as well as the form of the scalar field, which we will assume that inherits the spacetime symmetries and therefore is a function of the radial coordinate. Here $\alpha,\delta,\varepsilon,\zeta,\mu$ are constants of the theory. Note, that only derivatives of the scalar field appear in the action and hence we are free to substitute $\phi^{\prime}(r)$ with $X(r)$, using relation (10). Assuming the line element
@@
with $m$ being an integration constant related to the ADM mass of the black hole $\lim_{r\to\infty}(r-rf(r))/2=m$ . Note that the form of the kinetic energy does not contain the integration constant $m$ and that its behavior is solely determined by the constants of the theory. When $\alpha=\delta\zeta$, $X=0$ and f(r)=12m/rf(r)=1-2m/r, hence we go back to the Schwarzschild solution.
@@ III.1.2 Euclidean Action
where the subscript $E$ denotes Euclidean quantities, and $\psi$ collectively represents the matter fields. The spacetime element has an Euclidean signature $(+,+,+,+)$, and we can transition to this signature by performing a Wick rotation $t\to i\tau$. If we are considering a black hole spacetime, then to avoid the conical singularity at the horizon of the black hole $r_{h}$, the time coordinate $\tau$ must be periodic, with a period given by [121, 50]
@@
By imposing the saddle point approximation, we may consider that for the black hole spacetime that will concern us in this work, it is sufficient to consider the contribution of the action of the theory evaluated on the classical solution when the field equations hold: $\delta\mathcal{I}_{E}=0$ . Consequently, we need to calculate the on-shell action of our theory, ensuring it attains a true extremum within the class of fields considered. From this, we can relate the Euclidean action to the free energy of the system via
@@
$\mathcal{I}_{E}=\int d^{3}xd\tau\left[P\dot{\phi}+\Pi^{ij}\dot{g}_{ij}-NH-N^{i}H_{i}\right]+\mathcal{B}~,$ (21)
where $P$ is the corresponding conjugate momentum of the scalar field, $\Pi$ is the momentum of the metric, $N$ is the lapse function, $N^{i}$ is the shift vector, and $H,~H_{i}$ are the Hamiltonian densities. $\mathcal{B}$ denotes the boundary terms required to ensure a well-defined variational problem. For a very pedagogical explanation on why the boundary terms should be included, we refer the reader to [123]. Since both the scalar field and the spacetime metric are time independent, while the spacetime is non-rotating we can consider the reduced action principle
$\mathcal{I}_{E}=-4\pi\beta\int_{r_{h}}^{\infty}drNH+\mathcal{B}~,$ (22)
@@
This simple form of reduced Hamiltonian can be achieved by exploiting the fact that two theories are equivalent up to boundary terms. Hence when substituting the ADM metric (16) in the Euclidean action (21) we encounter terms like $\int dr(N^{\prime\prime}(r)\times A(r))$ where $A(r)$ is a function of $X,f$ and their derivatives. Such a term may be written as $\int dr(A^{\prime\prime}(r)\times N(r))$ where we have performed integration by parts and canceled the total derivatives that appeared. Consequently, we may rewrite the action (21) in terms of the dynamical fields $N,f,X$ as
$\mathcal{I}_{E}=4\pi\beta\int drN(r)\left(2rf^{\prime}(r)+2f(r)+\mu X(r)^{2}\left(\delta+\varepsilon r^{2}\right)+X(r)(\delta\zeta-\alpha)-2\right)+\mathcal{B}~.$ (24)
@@ III.1.3 Thermodynamic Quantities from Boundary Terms
$-\alpha+\delta\zeta+2\mu X(r)\left(\delta+\varepsilon r^{2}\right)=0~,$ (26)
@@
$\delta\mathcal{B}+8\pi\beta\int_{r_{h}}^{\infty}dr\frac{d}{dr}\left(r\delta f\right)=0~\to\delta\mathcal{B}=-8\pi\beta r\delta f(r)\big|_{r_{h}}^{\infty}~.$ (28)
@@
$\delta\mathcal{B}(\infty)=16\pi\beta\delta m~\hskip 8.5359pt\&~\hskip 8.5359pt\delta\mathcal{B}(r_{h})=-32\pi^{2}r_{h}\delta r_{h}=-4\pi\delta A(r_{h}),$ (31)
@@
Figure 1: The temperature as a function of the horizon radius. Here we have set $\delta\zeta/\alpha=4$$\delta/\varepsilon=2$ for three different ratios of $\mu/\zeta$.
@@ III.2 Parity Violating Theories
In this subsection we will discuss the thermodynamics of a black hole solution arising when the parity breaking term of $F_{5}(X)$ is present. As an example we will examine an analytic solution that has already been obtained in [27]. For $F_{4}(X)=0=G_{5}(X)$ and the coupling functions
@@
$\delta\mathcal{B}-8\pi\beta\gamma\int_{r_{h}}^{\infty}dr\frac{d}{dr}\left(\sqrt{f(r)}\delta f\right)=0~\to\delta\mathcal{B}=+8\pi\beta\gamma\sqrt{f(r)}\delta f(r)\big|_{r_{h}}^{\infty}~.$ (44)
@@
However, at the event horizon of the black hole we have $f(r_{h})=0$ and hence no boundary term is generated at the event horizon of the black hole, and by comparing with the previous discussion this will imply that these solutions have a constant entropy ($\delta\mathcal{S}=0$). Consequently, a well-defined variational principle implies